# Parallel Point-to-Point Shortest Paths and Batch Queries

### v1

@@ 7. Related Work
Existing Parallel Algorithms for BiDS and Batch PPSP.  Although BiDS is shown to be effective in accelerating sequential PPSP queries, we are unaware of any fully parallelized BiDS algorithm throughout the query phase. Some existing work (Rios and Chaimowicz, 2011; Vaira and Kurasova, 2011) can only parallelize bidirectional Dijkstra with two threads, one for the forward search and one for the backward search. PnP (Xu et al., 2019) uses bidirectional searches in preprocessing to predict the more efficient direction (from either the source or the target) to perform the search. Once predicted, it continues the search in only one direction, either forward or backward.

### v2

@@ 7. Related Work
Existing Parallel Algorithms for BiDS and Batch PPSP.  Although BiDS is shown to be effective in accelerating sequential PPSP queries, we are unaware of any fully parallelized BiDS algorithm throughout the query phase. Some existing works (Rios and Chaimowicz, 2011; Vaira and Kurasova, 2011) parallelize bidirectional Dijkstra with two threads, one for the forward search and one for the backward search, which naturally does not scale to more than two threads. PnP (Xu et al., 2019) uses bidirectional searches in preprocessing to predict the more efficient direction (from either the source or the target) to perform the search. Once predicted, it continues the search in only one direction, either forward or backward.
@@ Appendix C More Experimental Results
Large Batch Sizes on Batch PPSP.  We evaluate our batch PPSP algorithms on larger query batches in Fig. 8. We fix the number of query sources to 30 and increase the number of query pairs from 100 to 300. Overall, our Multi-BiDS algorithm consistently achieves the best geometric mean running time, outperforming the next best baseline by up to 1.32$\times$. As the number of query pairs increases, the performance gap between Multi-BiDS and the other algorithms becomes even more pronounced.
@@
Figure 8. Heatmap on large batch PPSP queries. The numbers are the running times normalized to the fastest algorithm on each test instance. Smaller or green is better. “GEOMEAN” $=$ geometric means on all graphs. We run two BiDS-based algorithms and two SSSP-based algorithms as introduced in Sec. 4. For BiDS-based algorithms, Multi is to evaluate all queries in a batch, Plain is to runs each query with our parallel BiDS algorithm one at a time, and Plain is to run all queries simultaneously in parallel, and each query is a parallel BiDS algorithm. For SSSP-based algorithms, VC is to run the SSSP algorithm from the minimum required sources induced by the vertex cover of the query graph, and Plain is to run the SSSP algorithm from all given sources.